# shapes.geometric, arrows

### v1

@@ 3 Preliminaries and Notation
General notation: We assume we are given a binary dataset $\mathcal{X}\in\{0,1\}^{N\times F}$ with $N$ samples, $F$ features, and associated binary labels $\mathcal{Y}\in\{0,1\}^{N}$. We let $[u]\coloneqq\{1,\ldots,u\}$, $I\subseteq[N]$ the indices of a subsample of the dataset, and $(x_{i},y_{i})$ denote the $i$th sample and its label. We define $\mathcal{X}|_{\mathcal{I}}\coloneqq\{x_{i}:i\in\mathcal{I}\}\subset\mathcal{X}$, $\mathcal{Y}|_{\mathcal{I}}\coloneqq\{y_{i}:i\in\mathcal{I}\}\subset\mathcal{Y}$, and $\mathcal{I}|_{f=k}\coloneqq\{i:i\in\mathcal{I}$ and $(x_{i})_{f}=k\}$, for $k\in\{0,1\}$. Finally, we let $c^{k}(\mathcal{I})$ be the count of points in $I$ with label $k\in\{0,1\}$, i.e., $c^{k}(\mathcal{I})=|\{i:i\in\mathcal{I}\text{ and }y_{i}=k\}|$ and $\mathcal{V}(I)$ be the set of nontrivial features splits of the samples in $I$, i.e., the set of features such that neither $I|_{f=0}$ nor $I|_{f=1}$ is nonempty.
@@ 3.1 AND/OR Graph Search
We briefly recapitulate the concept of AND/OR graphs and a search algorithm for AND/OR graphs, AO*. AND/OR graph search can be viewed as a generalization of the shortest path problem that allows nodes consisting of independent subproblems to be decomposed and solved separately. Thus, a solution of an AND/OR graph is not a path but rather a subgraph $\mathcal{S}$ with cost, denoted cost$(\mathcal{S})$, equal to the sum across the costs of its edges. AND/OR graphs contain two types of nodes: terminal nodes and nonterminal nodes. Nonterminal nodes can be further subdivided into AND nodes and OR nodes, with a special OR node designated as the root or start node $r$. A solution graph $\mathcal{S}$ on an AND/OR graph is connected subset of nodes of $\mathcal{G}$ in which:
@@
Figure 1: An example AND/OR graph, with AND nodes drawn as squares, and OR nodes drawn as solid circles, and terminal nodes drawn as dashed circles. The minimal cost solution is highlighted in red and has cost $0+0+3+4+1+2=10$.
@@
Additional AND/OR graph notation: In addition to the notation defined above, we use $t$ to refer to a terminal node. When searching over an AND/OR graph, we use $\mathcal{G}$ to refer to the implicit (entire) AND/OR graph and $\mathcal{G}^{\prime}\subset G$ to explicit (explored) AND/OR graph, as in prior work.
@@ 3.2 Bayesian Classification and Regression Trees (BCART)
We note that a BDT’s tree $T$ partitions the data such that the sample subsets $I(l_{1})$, $I(l_{2}),I(l_{L})$ fall into leaves $l_{1},l_{2},\dots,l_{L}$. Furthermore, a BDT defines a probability distribution over the respective labels occurring in their leaves: each label in leaf $l$ is sampled from Ber$(\theta_{l})$). Every BDT therefore induces a likelihood function, given in Theorem 1.
@@
$P(T|\mathcal{X})$ $=\left(\prod_{l\in T_{\text{leaves}}}p_{\text{leaf}}(d(l),I(l))\right)\times$
$\hskip 30.00005pt\left(\prod_{m\in T_{\text{internal}}}p_{\text{inner}}(d(m),I(m))\right)$
@@
$p_{\text{leaf}}(d,I)$ =$=\begin{cases}1,&\mathcal{V}(I)=\emptyset\\ 1-p_{\text{split}}(d),&\mathcal{V}(I)\neq\emptyset\end{cases}$ (6)
$p_{\text{inner}}(d,I)$ =$=\begin{cases}0,&\mathcal{V}(I)=\emptyset\\ p_{\text{split}}(d)/|\mathcal{V}(I)|,&\mathcal{V}(I)\neq\emptyset\end{cases}$ (7)
@@ 5 MAPTree
$\log\ell_{\text{leaf}}(c^{1}(I),c^{0}(I)),$ (11)
@@
$+\log\ell_{\text{leaf}}(c^{1}(I),0)$ (13)
$+\log\ell_{\text{leaf}}(0,c^{0}(I))\}$ (14)
and for AND node $a_{I,d,f}$ with OR node children $o_{I|_{f=0},d+1}$ and $o_{I|_{f=0},d+1}$, let
$h(a_{I,d,f})$ $=h(o_{I|_{f=0},d+1})+h(o_{I|_{f=0},d+1})$ (15)
@@ 5.1 Analysis of MAPTree
Algorithm 1 always outputs a minimal cost solution on $\mathcal{G}_{\mathcal{X},\mathcal{Y}}$ upon completion.
@@
Consider the tree induced by the output of Algorithm 1 under the natural bijection described in Section 4. By Theorems 5 and 6, this tree is the maximum a posteriori tree $\arg\max P(T|X,Y)$.
@@ 6.2 Fitting a Synthetic Dataset
In our experiments, MAPTree generates trees which outperform both the greedy, top-down approaches and ODT methods in test accuracy for various training dataset sizes and values of label corruption proportion $\epsilon$ (Figure 6). (We note that though some baseline algorithms demonstrate comparable performance at a single noise level, no baseline algorithm demonstrates test accuracy comparable to MAPTree across all noise levels). We also emphasize that MAPTree requires no hyperparameter tuning, whereas we experimented with various values of hyperparameters for the baseline algorithms in which performance was highly dependent on hyperparameter values (e.g., DL8.5 and GOSDT); see Appendix B.
@@ Appendix A Proofs of Theorems
We prove this via induction on iteration. After the first iteration, there is only terminal node in $\mathcal{G}^{\prime}$, and only one valid solution $\mathcal{S}_{0}$ exists in $\mathcal{G}^{\prime}$: $\mathcal{S}_{0}=\{o_{[N],0},t_{[N],0}\}$. Thus, no nodes other than $r=o_{[N],0}$ have valid partial solutions. At this point, $UB[r]=\texttt{cost}(o_{[N],0},t_{[N],0})=\texttt{cost}(\mathcal{S})$ and $UB$ is undefined on all other nodes, as required.
@@
When $|\mathcal{I}|\leq 1$, any split will lead to an empty subtree, meaning the cost of splitting is always $\infty$, and the minimal cost of any partial solution rooted at $o_{\mathcal{I},d}$ is either $\texttt{cost}(o_{\mathcal{I},d},t_{\mathcal{I},d})$ or +$\infty$+ ∞ if $t_{\mathcal{I},d}\not\in\mathcal{G}^{\prime}$. If $t_{\mathcal{I},d}\not\in\mathcal{G}^{\prime}$, this implies that $o_{\mathcal{I},d}\not\in\mathcal{E}$, meaning $UB[o_{\mathcal{I},d}]$ is undefined. If $t_{\mathcal{I},d}\in\mathcal{G}^{\prime}$ and $|\mathcal{I}|\leq 1$, then $UB[o_{\mathcal{I},d}]$ is $\texttt{cost}(o_{\mathcal{I},d},t_{\mathcal{I},d})$, as required.
@@
We will show that getSolution($o_{\mathcal{I},d}$) outputs a minimal cost partial solution of $\mathcal{G}^{\prime}$ rooted at OR node $o_{\mathcal{I},d}$ via induction on $|\mathcal{I}|$. When $|\mathcal{I}|\leq 1$, any splits will lead to a empty subtrees, meaning the cost of splitting is always $\infty$, so $\texttt{getSolution}(o_{\mathcal{I},d})$ will return $\{o_{\mathcal{I},d},t_{\mathcal{I},d}\}$. When $|\mathcal{I}|>1$, getSolution($o_{\mathcal{I},d}$) will either stop for a minimal cost or recurse on a split that yields minimal $UB$ value. Lemma 16 shows that the $UB$ values of $o_{\mathcal{I},d}$ and its children are equal to the minimal cost across all partial solutions rooted at each of these respective nodes. As a result, if getSolution stops, then $\{o_{\mathcal{I},d},t_{\mathcal{I},d}\}$ is a minimal cost partial solution. Otherwise, if getSolution splits on feature $f$, $\{o_{\mathcal{I},d},a_{\mathcal{I},d,f}\}\,\cup\texttt{getSolution}(o_{\mathcal{I}|_{f=0},d+1})\,\cup\texttt{getSolution}(o_{\mathcal{I}|_{f=0},d+1})$ is also a minimal cost partial solution by the inductive hypothesis. ∎
@@
The lower bounds $LB$ represent correct lower bounds on the true value of a node in every iteration. For any OR node $o$, we have that $LB[o]\leq f(o)$ and for any AND node $a$, we have that $LB[a]\leq f(a)$.
@@
We prove the contrapositive. Consider any OR node $o_{\mathcal{I},d}$. We show that if every OR node descendant of $o_{\mathcal{I},d}$ is in $\mathcal{E}$, then $LB[o_{\mathcal{I},d}]=UB[o_{\mathcal{I},d}]$. We show this via induction on $|\mathcal{I}|$. When $|\mathcal{I}|\leq 1$, splitting further incurs infinite cost, so $LB[o_{\mathcal{I},d}]=UB[o_{\mathcal{I},d}]=\texttt{cost}(o_{\mathcal{I},d},t_{\mathcal{I},d})$. When $|\mathcal{I}|>1$, if every OR node descendant of $o_{\mathcal{I},d}$ is in $\mathcal{E}$, then $o_{\mathcal{I},d}\in\mathcal{E}$ must also hold. Since the OR node descendants of the children of $o_{\mathcal{I},d}$ must be in $\mathcal{E}$ as well, they must all have matching $UB$ and $LB$ by inductive hypothesis, meaning $LB[o_{\mathcal{I},d}]=UB[o_{\mathcal{I},d}]$. ∎

### v3

# MAPTree: Beating “Optimal” Decision Trees with Bayesian Decision Trees

@@ 3 Preliminaries and Notation
General notation: We assume we are given a binary dataset $\mathcal{X}\in\{0,1\}^{N\times F}$ with $N$ samples, $F$ features, and associated binary labels $\mathcal{Y}\in\{0,1\}^{N}$. We let $[u]\coloneqq\{1,\ldots,u\}$, $\mathcal{I}\subseteq[N]$ the indices of a subsample of the dataset, and $(x_{i},y_{i})$ denote the $i$th sample and its label. We define $\mathcal{X}|_{\mathcal{I}}\coloneqq\{x_{i}:i\in\mathcal{I}\}\subset\mathcal{X}$, $\mathcal{Y}|_{\mathcal{I}}\coloneqq\{y_{i}:i\in\mathcal{I}\}\subset\mathcal{Y}$, and $\mathcal{I}|_{f=k}\coloneqq\{i:i\in\mathcal{I}$ and $(x_{i})_{f}=k\}$, for $k\in\{0,1\}$. Finally, we let $c^{k}(\mathcal{I})$ be the count of points in $\mathcal{I}$ with label $k\in\{0,1\}$, i.e., $c^{k}(\mathcal{I})=|\{i:i\in\mathcal{I}\text{ and }y_{i}=k\}|$ and $\mathcal{V}(\mathcal{I})$ be the set of nontrivial features splits of the samples in $\mathcal{I}$, i.e., the set of features such that neither $\mathcal{I}|_{f=0}$ nor $\mathcal{I}|_{f=1}$ is nonempty.
@@ 3.1 AND/OR Graph Search
We briefly recapitulate the concept of AND/OR graphs and a search algorithm for AND/OR graphs, AO*. AND/OR graph search can be viewed as a generalization of the shortest path problem that allows nodes consisting of independent subproblems to be decomposed and solved separately. Thus, a solution of an AND/OR graph is not a path but rather a subgraph $\mathcal{S}$ with cost, denoted cost$(\mathcal{S})$, equal to the sum across the costs of its edges. AND/OR graphs contain two types of nodes: terminal nodes and nonterminal nodes. Nonterminal nodes can be further subdivided into AND nodes and OR nodes, with a special OR node designated as the root or start node $r$. For a given AND/OR graph $\mathcal{G}$, a solution graph $\mathcal{S}$ on an AND/OR graph is a connected subset of nodes of $\mathcal{G}$ in which:
@@
Figure 1: An example (general) AND/OR graph, with AND nodes drawn as squares, and OR nodes drawn as solid circles, and terminal nodes drawn as dashed circles. The minimal cost solution is highlighted in red and has cost $0+0+3+4+1+2=10$. This diagram demonstrates an AND/OR graph where the root node $r$ is an AND node; in MAPTree, the root node is an OR node.
@@
Additional AND/OR graph notation: In addition to the notation defined above, we use $t$ to refer to a terminal node. When searching over an AND/OR graph, we use $\mathcal{G}$ to refer to the implicit (entire) AND/OR graph and $\mathcal{G}^{\prime}\subset\mathcal{G}$ to explicit (explored) AND/OR graph, as in prior work.
@@ 3.2 Bayesian Classification and Regression Trees (BCART)
We note that a BDT’s tree $T$ partitions the data such that the sample subsets $\mathcal{I}(l_{1})$, $\mathcal{I}(l_{2}),\mathcal{I}(l_{L})$ fall into leaves $l_{1},l_{2},\dots,l_{L}$. Furthermore, a BDT defines a probability distribution over the respective labels occurring in their leaves: each label in leaf $l$ is sampled from Ber$(\theta_{l})$). Every BDT therefore induces a likelihood function, given in Theorem 1.
@@
$P(T|\mathcal{X})$ $=\left(\prod_{l\in T_{\text{leaves}}}p_{\text{leaf}}(d(l),\mathcal{I}(l))\right)\times$
$\hskip 30.00005pt\left(\prod_{m\in T_{\text{internal}}}p_{\text{inner}}(d(m),\mathcal{I}(m))\right)$
@@
$p_{\text{leaf}}(d,\mathcal{I})$ =$=\begin{cases}1,&\mathcal{V}(\mathcal{I})=\emptyset\\ 1-p_{\text{split}}(d),&\mathcal{V}(\mathcal{I})\neq\emptyset\end{cases}$ (6)
$p_{\text{inner}}(d,\mathcal{I})$ =$=\begin{cases}0,&\mathcal{V}(\mathcal{I})=\emptyset\\ p_{\text{split}}(d)/|\mathcal{V}(\mathcal{I})|,&\mathcal{V}(\mathcal{I})\neq\emptyset\end{cases}$ (7)
@@ 5 MAPTree
$\log\ell_{\text{leaf}}(c^{1}(\mathcal{I}),c^{0}(\mathcal{I})),$ (11)
@@
$+\log\ell_{\text{leaf}}(c^{1}(\mathcal{I}),0)$ (13)
$+\log\ell_{\text{leaf}}(0,c^{0}(\mathcal{I}))\}$ (14)
and for AND node $a_{\mathcal{I},d,f}$ with OR node children $o_{\mathcal{I}|_{f=0},d+1}$ and $o_{\mathcal{I}|_{f=1},d+1}$, let
$h(a_{\mathcal{I},d,f})$ $=h(o_{\mathcal{I}|_{f=0},d+1})+h(o_{\mathcal{I}|_{f=1},d+1})$ (15)
@@ 5.1 Analysis of MAPTree
When Algorithm 1 does not terminate early due to the time remaining condition, it always outputs a minimal cost solution on $\mathcal{G}_{\mathcal{X},\mathcal{Y}}$ upon completion.
@@
Consider the tree induced by the output of Algorithm 1 under the natural bijection described in Section 4. By Theorems 5 and 6, this tree is the maximum a posteriori tree $\arg\max_{T}P(T|\mathcal{X},\mathcal{Y})$.
@@ 6.2 Fitting a Synthetic Dataset
In our experiments, MAPTree generates trees which outperform both the greedy, top-down approaches and ODT methods in test accuracy for various training dataset sizes and values of label corruption proportion $\epsilon$; the results are presented in Figure 6 in Appendix B due to space constraints. We note that though some baseline algorithms demonstrate comparable performance at a single noise level, no baseline algorithm demonstrates test accuracy comparable to MAPTree across all noise levels. We also emphasize that MAPTree requires no hyperparameter tuning, whereas we experimented with various values of hyperparameters for the baseline algorithms in which performance was highly dependent on hyperparameter values (e.g., DL8.5 and GOSDT); see Appendix B.
@@ 7 Discussion and Conclusions
Acknowledgements
We would like to thank the anonymous reviewers and Area Chair for their reviews and helpful feedback.
M. T. was funded by a J.P. Morgan AI Fellowship, a Stanford Indisciplinary Graduate Fellowship, a Stanford Data Science Scholarship, and an Oak Ridge Institute for Science and Engineering Fellowship.
@@ Appendix A Proofs of Theorems
We prove this via induction on iteration. After the first iteration, there is only terminal node in $\mathcal{G}^{\prime}$, and only one valid solution $\mathcal{S}$ exists in $\mathcal{G}^{\prime}$: $\mathcal{S}_{0}=\{o_{[N],0},t_{[N],0}\}$. Thus, no nodes other than $r=o_{[N],0}$ have valid partial solutions. At this point, $UB[r]=\texttt{cost}(o_{[N],0},t_{[N],0})=\texttt{cost}(\mathcal{S})$ and $UB$ is undefined on all other nodes, as required.
@@
When $|\mathcal{I}|=1$, any split will lead to an empty subtree, meaning if $t_{\mathcal{I},d}\in\mathcal{G}^{\prime}$ then $\{o_{\mathcal{I},d},t_{\mathcal{I},d}\}$ is the only partial solution rooted at $o_{\mathcal{I},d}$ and otherwise no such partial solution exists. If $t_{\mathcal{I},d}\not\in\mathcal{G}^{\prime}$, this implies that $o_{\mathcal{I},d}\not\in\mathcal{E}$, meaning $UB[o_{\mathcal{I},d}]$ is undefined. If $t_{\mathcal{I},d}\in\mathcal{G}^{\prime}$ and $|\mathcal{I}|=1$, then $UB[o_{\mathcal{I},d}]$ is $\texttt{cost}(o_{\mathcal{I},d},t_{\mathcal{I},d})$, as required.
@@
We will show that getSolution($o_{\mathcal{I},d}$) outputs a minimal cost partial solution of $\mathcal{G}^{\prime}$ rooted at OR node $o_{\mathcal{I},d}$ via induction on $|\mathcal{I}|$. When $|\mathcal{I}|=1$, any splits will lead to a empty subtrees, so $\texttt{getSolution}(o_{\mathcal{I},d})$ must return $\{o_{\mathcal{I},d},t_{\mathcal{I},d}\}$. When $|\mathcal{I}|>1$, getSolution($o_{\mathcal{I},d}$) will either stop for a minimal cost or recurse on a split that yields minimal $UB$ value. Lemma 16 shows that the $UB$ values of $o_{\mathcal{I},d}$ and its children are equal to the minimal cost across all partial solutions rooted at each of these respective nodes. As a result, if getSolution stops, then $\{o_{\mathcal{I},d},t_{\mathcal{I},d}\}$ is a minimal cost partial solution. Otherwise, if getSolution splits on feature $f$, $\{o_{\mathcal{I},d},a_{\mathcal{I},d,f}\}\,\cup\texttt{getSolution}(o_{\mathcal{I}|_{f=0},d+1})\,\cup\texttt{getSolution}(o_{\mathcal{I}|_{f=0},d+1})$ is also a minimal cost partial solution by the inductive hypothesis. ∎
@@
The lower bounds $LB$ represent correct lower bounds on the true value of a node in every iteration. For any OR node $o$ with true value $f(o)$, we have that $LB[o]\leq f(o)$ and for any AND node $a$ with true value $f(a)$, we have that $LB[a]\leq f(a)$.
@@
We prove the contrapositive. Consider any OR node $o_{\mathcal{I},d}$. We show that if every OR node descendant of $o_{\mathcal{I},d}$ is in $\mathcal{E}$, then $LB[o_{\mathcal{I},d}]=UB[o_{\mathcal{I},d}]$. We show this via induction on $|\mathcal{I}|$. When $|\mathcal{I}|=1$, splitting further incurs infinite cost, so $LB[o_{\mathcal{I},d}]=UB[o_{\mathcal{I},d}]=\texttt{cost}(o_{\mathcal{I},d},t_{\mathcal{I},d})$. When $|\mathcal{I}|>1$, if every OR node descendant of $o_{\mathcal{I},d}$ is in $\mathcal{E}$, then $o_{\mathcal{I},d}\in\mathcal{E}$ must also hold. Since the OR node descendants of the children of $o_{\mathcal{I},d}$ must be in $\mathcal{E}$ as well, they must all have matching $UB$ and $LB$ by inductive hypothesis, meaning $LB[o_{\mathcal{I},d}]=UB[o_{\mathcal{I},d}]$. ∎
@@ B.2 Additional Experiments
In this subsection, we present additional experimental results that were omitted from the main paper due to space constraints.
Figure 6 shows that MAPTree generates trees which out-perform both the greedy, top-down approaches and ODT methods in test accuracy for various training dataset sizes and values of label corruption proportion $\epsilon$.